# The Role of Human Creativity in the Presence of AI Creativity Tools at Work: A Case Study on AI-Driven Content Transformation in Journalism

### v1

@@ Abstract.
As AI becomes more capable, it is unclear how human creativity will remain essential in jobs that incorporate AI. We conducted a 14-week study of a student newsroom using an AI tool to convert web articles into social media videos. Most treated the tool as a creative springboard, yet still had to edit many AI outputs. The tool enabled the team to publish successful content, receiving over 500,000 views. Yet creators sometimes treated AI as an unquestioned expert, accepting flawed suggestions. Editorial critique was essential to spot errors and guide creative solutions when AI failed. We discuss how AI’s inherent gaps ensure human creativity remains vital.
@@ 1. Introduction
The integration of AI into creative processes is revolutionizing content creation, particularly in the fast-paced environment of newsrooms. As generative AI tools become increasingly prevalent, understanding the intricate interactions between these technologies and human creativity is crucial. What is the role of human creativity in the presence of AI creativity tools in the workplace? Does the role of human creativity change over time as people gain familiarity with AI tools? Despite the growing adoption of AI in creative domains, these questions remain underexplored. Although AI has some capacity for creation, creativity is a complex cognitive process involving many skills, including convergent thinking, divergent thinking, framing, analysis, critique, audience understanding, and iteration. AI may be able to assist with some or parts of these processes, but human creativity is likely to still play a large role in the full creative process.
@@
To explore the role of human creativity in the presence of AI, this paper presents a 14-week study conducted within the social media team of a university journalism department’s newsroom, focusing on content creation for the TikTok platform. The team included five social media creators (all students), one editor (a faculty member), and two market researchers (research assistants), embodying a typical real-world newsroom structure. They used an AI tool designed to help creators transform web articles into engaging short video scripts and storyboards. Using the tool, the journalists read, selected, regenerated, and edited several possible AI-generated scripts before shooting and editing the videos themselves. Each journalist aimed to produce one new video a week and iterate on previous videos. The newsroom met weekly for the editor to review videos, request changes, and decide if a video was ready to publish. After 10 weeks, the market researchers conducted 42 interviews with members of the target audience to assess quality and gather suggestions for improvement. The study is complemented by comparative insights from a shorter case study involving two professional TV journalists.
We found that overall, creators used AI as a “creative springboard.” Four out of five creators consistently used the tool and they appreciated its ability to reframe news stories around characters—part of the novel social media format. Without the tool, it was difficult to balance information and entertainment. However, for final scripts, they often edited “about half” of the text for accuracy and clarity. Creators did choose to inject their own creative touches, especially in areas the tool did not support, such as adding visual elements like props or animations or using their acting ability to bring new emotions, attitudes, and depth to their characters. Creators sometimes chose not to use the tool when they felt confident about the story or had a specific creative inspiration. Overall, the team was successful in publishing 16 videos in 14 weeks, and amassing over 500,000 views.
@@
Figure 1. An example of a news reel made during the newsroom study describes the three key facts of a story about a teacher shortage.
[…]
An example of a news reel made during the newsroom study describes the three key facts of a story about a teacher shortage. There are eight frames with two characters: a school board and a teachers union. thru dialog, they describe the problems the district faces with teacher shortages.
Human critical evaluation was essential in defending against AI errors. Surprisingly, creators often accepted poor AI suggestions, such as overly positive language and meaningless phrases. This was true of both student creators and experienced, professional TV journalists. Although journalists are experts in journalism, they are not experts in the new skill of making videos. Here, AI is perceived as an expert. Journalists defer to AI suggestions, which are coherent even if flawed in other ways. Additionally, AI can fail dramatically when its lack of context results in socially inappropriate suggestions. These flaws required creators’ critical intervention either to modify these narratives or to discard them entirely to uphold social appropriateness.
Creators’ relationship with the AI changed over the course of the 14-week study. Initially, they relied heavily on the AI’s outputs, but with time, creators demonstrated a growing tendency to integrate their creativity outside the AI’s direct outputs. When the team gathered feedback from target viewers, they found even more ways to make their content more engaging, beyond what the tool was offering. Gradual adaptation, editorial interventions, and learning from audience feedback over time all encouraged creators to continue to use their own creativity on top of the creativity support the tool offered.
@@ 2. Background and Related Work
Figure 2. ReelFramer is a human-AI co-creative system that supports journalists in creating news reels by transforming web articles into scripts and storyboards. ReelFramer enables the exploration of multiple narrative framings that span the infotainment spectrum and their foundational details. It also aids in visual framing for character and key visual detail design.
@@ 2.1. Studies of AI in a Work Context
Recent studies have mapped out the significant opportunities that generative AI brings to various industries, highlighting its potential to enhance productivity for a wide range of workspace tasks (Brynjolfsson et al., 2023; Noy and Zhang, 2023). It has shown promise in replacing both cognitive routine work (often considered drudgery) and cognitive non-routine work, including creative tasks (Bouschery et al., 2023). In an experimental study with a writing task carried out by 453 college-educated professionals, Noy and Zhang (2023) found that those using ChatGPT both saved time and increased quality. Two experiments with Boston Consulting Group (Dell’Acqua et al., 2023) show that consultants with access to ChatGPT were more productive and produced higher-quality ideas. Studies of AI coding tools show it leads to a 26% increase in weekly coding tasks completed over three companies (Cui et al., 2024; Peng et al., 2023).
However, various studies also express reservations about AI’s ability to fully replace human workers. A study of creativity support tools demonstrates that while the tools are helpful for specific tasks, they cannot solve problems entirely on their own (Tao et al., 2023)—especially for tasks involving core human skills, even in the long run (Long et al., 2024). AI support can even be problematic. Dell’Acqua et al. (2023) show that for difficult tasks, subjects using ChatGPT were about 20% less likely to make the correct recommendation, as many were led astray by ChatGPT’s convincing but incorrect analysis. Even when AI is correct, users face challenges incorporating AI results (Kim and Kang, 2024).
These limitations suggest that AI might not completely reshape entire industries but rather impact certain aspects of how people work and what they focus on (Randazzo et al., 2024). However, developing these new configurations requires learning which tasks AI can be trusted to complete and how best to incorporate its results.
@@ 2.2. AI for Creativity Support
The use of AI for creativity support, particularly to assist writing, has expanded rapidly in recent years, with applications across various domains, including journalism (Liu et al., 2022; Wang et al., 2024a; Petridis et al., 2023). AI writing assistants can provide scaffolding for writers by supporting different cognitive stages of the writing process (Flower, 1981), including framing (Hui and Sprouse, 2023; Wang et al., 2024a), idea generation (Clark et al., 2018; Petridis et al., 2023) to drafting (Gero et al., 2022; Zhang et al., 2023; Long et al., 2023) and revision (Lee et al., 2024; Gero et al., 2022; Wu et al., 2019). Lee et al. (Lee et al., 2024) provide a comprehensive design space for intelligent and interactive writing assistants that emphasizes the importance of considering task, user, technology, interaction, and ecosystem aspects when using AI to support writers. However, using AI in creativity also has risks; overreliance (Ashktorab et al., 2021; Kim et al., 2024; Vasconcelos et al., 2023; Zhang et al., 2024) on AI can cause people to accept poor AI answers, and AI ideation can actually reduce diversity across people (Anderson et al., 2024). Broadly, AI can be seen not as a solution, but as a design material that needs to be understood by human creators (Yildirim et al., 2022; Subramonyam et al., 2021).
A major creative application of AI lies in content retargeting and transformation—editing and repurposing text, image, video, or audio for new audiences, platforms, or goals. The key is to find the fundamental message to preserve while adapting the content into the new format. Previous works have explored how to support people with content transformation tasks in terms of video (Truong et al., 2021; Chi et al., 2021, 2022; Wang et al., 2024b, 2022). For example, AI systems can generate video podcast trailers by identifying salient scenes to summarize content while preserving its tone (Wang et al., 2024b), or shorten audio stories for social media consumption (Wang et al., 2022). However, creating effective short-form content poses unique challenges, particularly in planning and structuring narratives (Kim and Kim, 2024). Systems like ReelFramer (Wang et al., 2024a) address these hurdles by embedding genre-specific knowledge to guide writers in scripting videos that align with journalistic formats, exemplifying how AI can scaffold human creativity in multimedia journalism. By encoding domain expertise, such tools streamline the adaptation of long-form content into short, platform-optimized outputs.
@@ 2.3. News Reels
News reels help outlets reach younger viewers who favor TikTok and Instagram (Newman, 2022). Nearly half of major publishers now post on TikTok, and younger audiences are increasingly seeking news there (Newman et al., 2022). Although both news reels and traditional news aim to educate the audience on relevant information, the narrative forms and framings are very different. Traditional journalistic narratives are fact-driven and explain events in a serious tone. By contrast, news reels seek to balance information and entertainment by taking a more playful tone, in order to conform to the narrative structures common to social media videos. A common way of achieving this goal is to act out the story from the perspective of different characters that personify aspects of the story. For example, in a Washington Post reel [^1] https://www.tiktok.com/@washingtonpost/video/7147811605655473450, a character representing Hurricane Ian explains to another character representing the hurricane’s roommate why he is “visiting” Florida and all the trouble he might cause there. Even for a serious topic like hurricane destruction, news reels must strike a balance between informational content and entertainment value (Newman, 2022). Figure 1 contains an example of a news reel made in our newsroom study with a dialog between a School District character and a Teachers Union character acting out the news of a teacher shortage. As is typical of news reels, the creator is playing both roles.
Even within the role-play format, framing a news story for social media is challenging. In interviews conducted by the Reuters Institute in 2022 with news reel creators (Newman, 2022), many different approaches to creating reels have been found, but a key challenge is to balance between information and entertainment (Newman, 2022). If a serious topic is portrayed too light-heartedly, it may diminish the perceived seriousness of the issue (Davis et al., 2022). However, if a video does not grab a reader’s attention with something that is interesting and entertaining, they may scroll past it. Some news organizations take a conservative approach and stay closer to the TV journalism format, but others lean into the tropes of social media videos to blend news information into an engaging social media style. The challenge we address is to help journalists translate their articles into this style, skillfully blending information and entertainment to strike an appropriate tone.
@@ 2.4. ReelFramer AI Tool
ReelFramer (Wang et al., 2024a) is a creativity support system that uses generative text and image AI to scaffold the process of creating a script and storyboard for a news reel. Generative AI is helpful for exploring multiple narrative framings and their details, and generating scripts, character boards, and storyboards. The system is designed to be highly interactive, allowing people to make choices and guide the system at every step to ensure the news reels it produces are accurate, authentic, and achievable to execute.
The user starts by inputting the news article (see Figure 2). The system uses a large language model (LLM), GPT-4, to extract locations, people, and activities from the article. The user then selects one of the three narrative framings, expository dialogue, reenactment, or comedic analogy. Given a framing, the system uses the LLM to suggest foundational details like characters, plot, setting, and key information, which is also called premise in scriptwriting (Batty et al., 2017). The user can accept, regenerate, or edit these suggestions. Once the user finalizes a premise, the system automatically generates a script. Again, the user can accept, regenerate, or edit the script. Users are encouraged to explore many narrative framings, premises, and scripts. Once the user is happy with the script, the system provides visual framing to explore the characters’ visual design. It generates a character board that contains visual details of the costumes, props, and backgrounds to distinguish the characters using a text-to-image model, DALL-E 2. Based on the character board, it generates a storyboard that shows suggested emotions, actions, and dialogues for each character. Users can then make reels based on these materials.
3. Methods
3.1. Overview
We conducted three studies: 1) a 14-week study of how a small newsroom adopts a creativity support tool, 2) an audience reception study assessing the videos made with the tool, and 3) a 4-week study of how two professional TV journalists adopt a creativity support tool. We used iterative, exploratory methods (Creswell and Clark, 2017), guided by early insights from student journalists. Each study strand was analyzed independently yet informed by prior findings. Institutional review boards deemed the project exempt.
3.2. Organizational Cohort Field Study with Student Journalists
In the first study, researchers recruited advanced students from a university journalism program that runs a TV and print journalism media outlet called AnonymizedNews to be part of researchers’ simulated newsroom. As part of their existing program, the students report local area news, often publishing multiple stories a day with faculty guidance and approval. AnonymizedNews  has a fledgling TikTok channel, with fewer than 50 posts and no posts getting more than 3000 views. Simulated newsroom environments are common in journalism education spaces (Steel et al., 2007) as part of practical and experiential learning with both procedural (environment) and factual (task) authenticity (Chen et al., 2001). Choosing students in a university newsroom was fairly representative of professional newsrooms because they often hire recent graduates to run social media, due to their familiarity with social media and the demographic is aspired to reach.
@@ 3.2. Organizational Cohort Field Study with Student Journalists
In this simulated newsroom environment, we had one researcher acting as editor, two researchers acting as market researchers, and five students given news stories each week to convert to reels, as if they were part of a digital news desk. The news stories were ones the newsroom had previously reported, sometimes by the student themselves. In weekly meetings, researchers acted as both participant observers  (Emerson et al., 2011), taking notes on how student journalists used the tool, and observant participants, interacting with the student journalists to offer the support an editor would provide (Seim, 2024), while allowing space for participants to share insights on production techniques with fellow student journalists as they would in a standard operating newsroom. Student journalists were regularly asked about their use or nonuse of the tool in the creation of their newsreels. The editor led group feedback discussions for every video, and gave specific feedback for iteration. After approval by the editor, newsreels were published to the AnonymizedNews  TikTok channel.
3.3. Audience Reception Study
To compare the quality of the student journalist-produced news reels to the other short form video news content (Shearer, 2024), we conducted qualitative interviews with typical short-form video audience members (Brennen, 2021). This allowed us to do two things: 1) access the relative quality of the student production, and 2) provide the student journalists with specific information to improve the audience reception of their videos. Questions covered how informative and entertaining the video was, their likelihood to watch the video if they came across it on their personal feed, and the likelihood of the video fitting in with content they typically see, from both qualitative and quantitative perspectives. Interviews took between 20 and 66 minutes, with an average of 35 minutes. 46 individuals signed up to participate, n=42 participants were interviewed.
Participants were shown six total short-form news videos. All six videos were systematically shown in a predetermined balanced design to evenly distribute order effects. Participants watched four videos created by the student journalists using ReelFramer on a range of topics. Two additional videos were added for comparison: one was professionally created by the Washington Post’s TikTok team to cover the trajectory of a hurricane, the other was created by a student in the AnonymizedNews  program but not in the study and not using the tool; it discussed a local basketball tournament.
Participants agreed to be interviewed in exchange for course credit as part of the university’s subject pool system. All participants were university students from across disciplines aged between 18-30, ($\mu$ = 20). Researchers recorded the interviews over Zoom and transcribed the videos before analysis began for axial themes (Saldana, 2003). The distribution of quantitative responses was plotted to compare the informativeness and fit of the news reels.
3.4. Qualitative Cohort Field Study with Practicing Journalists
To understand if there were differences in how professional journalists would adapt to the creativity support tool, researchers conducted a qualitative cohort field study with practicing journalists. In the summer of 2024, researchers recruited a cohort of journalists by posting in a social media group for multimedia journalists. Researchers informed participants we would hold four, one-hour collaborative sessions. 12 individuals completed a demographic questionnaire. Four individuals attended at least one training session, two produced at least one video using the tool. This sample size and these levels of attrition are not unexpected in longitudinal studies and do not diminish the benefit of the findings (Saldana, 2003). Similar to the student study, researchers acted in the capacity of newsroom editor offering critique and advice.
4. Study 1 Findings: Newsroom Field Study
4.1. AI Usage: AI as a Creative Springboard
@@
Overall, the student group made 26 news reels, 16 of which were published by the editor. The 10 that were unpublished were still being iterated at the end of the session or had been abandoned (Section 4.2.2). Altogether, the 16 videos published by the team garnered over 500,000 views. Four of the five student journalists chose to use the tool for all or most of their videos. 13 of 16 published videos used ReelFramer in a significant manner to help achieve an output the editor deemed publishable.
4.1.1. Reasons for AI Tool Usage
Creators choose to use the AI tool a majority of the time—for 20 of 26 reels created. Creators who chose to use the tool often described it as “a creative springboard.” There were two main uses.

### v2

@@ Abstract.
As AI becomes more capable, it is unclear how human creativity will remain essential in jobs that incorporate AI. We conducted a 14-week study of a student newsroom using an AI tool to convert web articles into social media videos. Most creators treated the tool as a creative springboard, not as a completion mechanism. They edited the AI outputs. The tool enabled the team to publish successful content that received over 500,000 views. Human creativity remained essential: after AI produced templated outputs, creators took ownership of the task, injecting their own creativity, especially when AI failed to create appropriate content. AI was initially seen as an authority, due to creators’ lack of experience, but they ultimately learned to assert their own authority.
@@ 1. Introduction
The integration of AI into creative processes is revolutionizing content creation, particularly in the fast-paced environment of newsrooms. As generative AI tools become increasingly prevalent, understanding the intricate interactions between these technologies and human creativity is crucial. What is the role of human creativity in the presence of AI creativity tools in the workplace? Does the role of human creativity change over time as people gain familiarity with AI tools? Although AI has some capacity for creation, creativity is a complex cognitive process involving many skills, including divergent and convergent thinking, framing, analysis, critique, audience understanding, and iteration. AI may be able to assist with some or parts of these processes, but human creativity is likely to still play a role in the creative process.
@@
To explore the role of human creativity in the presence of AI, this paper presents a 14-week study conducted within the social media team of a university journalism department’s newsroom, focusing on content creation for the TikTok platform. The team included five social media creators (all students), one editor (a faculty member), and two market researchers (research assistants), embodying a typical real-world newsroom structure. They used an AI tool designed to help creators transform web articles into engaging short video scripts and storyboards. Using the tool, the journalists read, selected, regenerated, and edited several possible AI-generated scripts before shooting and editing the videos themselves. Each journalist aimed to produce one new video a week and iterate on previous videos. The newsroom met weekly for the editor to review videos, request changes, and decide if a video was ready to publish. After 10 weeks, the market researchers conducted 42 interviews with members of the target audience to gather suggestions for improvement.
@@
Figure 1. A news reel made during the student study covering local news about a teacher shortage.
Overall, the team was successful in publishing 16 videos in 14 weeks and amassing over 500,000 views. AI was far from perfect—creators had to edit “about half” of the AI scripts for accuracy and clarity. However, AI was helpful both as scaffolding when learning the task, and as a creative springboard (Shneiderman, 2007) even after they understood the task. Additionally, the tool was not strictly necessary—creators opted not to use the tool when it was not working well for a story or if they had a specific creative inspiration. In addition to correcting AI, creators also intervened to inject their own creativity in many parts of the process. Whereas AI could only produce outputs, the creators took ownership over the task, using their judgment and creativity to maximize its value to audiences, rather than just adapting a story to a template as AI did. The creators’ relationship with AI changed over the course of the 14-week study. Initially, they relied heavily on the AI’s outputs—even accepting its blatant errors without question. However, through critical reflection and iteration, creators were able to reassert themselves. AI was initially seen as “an authority” on news reels, due to creators’ lack of experience and AI’s mystique of intelligence. But ultimately, users were better able to balance AI’s authority and their own.
@@ 2. Background and Related Work
Figure 2. ReelFramer is a human-AI co-creative system that supports journalists in creating news reels from print articles. Users input the text of a news article, and accept or reject AI suggestions for news information, narrative framing elements, script, visual framing, and a storyboard.
@@ 2.1. Studies of AI in a Work Context
Recent studies have mapped out the significant opportunities that generative AI brings to various industries, highlighting its potential to enhance productivity for a wide range of workspace tasks (Brynjolfsson et al., 2023; Noy and Zhang, 2023). It has shown promise in replacing both cognitive routine work (often considered drudgery) and cognitive non-routine work, including creative tasks (Bouschery et al., 2023), professional writing tasks (Noy and Zhang, 2023), professional software development (Cui et al., 2024; Peng et al., 2023), and professional analysis work (Dell’Acqua et al., 2023).
However, various studies also express reservations about AI’s ability to fully replace human workers. A study of creativity support tools demonstrates that while the tools are helpful for specific tasks, they cannot solve problems entirely on their own (Tao et al., 2023)—especially for tasks involving core human skills, even in the long run (Long et al., 2024). AI support can even be problematic. (Dell’Acqua et al., 2023) shows that for difficult tasks, subjects using ChatGPT were about 20% less likely to make the correct recommendation, as many were led astray by its convincing but incorrect analysis. Even when AI is correct, users face challenges incorporating AI results (Kim and Kang, 2024). These limitations suggest that AI might not completely reshape entire industries but rather impact certain aspects of how people work and what they focus on (Randazzo et al., 2024).
@@ 2.2. AI for Creativity Support
The use of AI for creativity support, particularly to assist writing, has expanded rapidly in recent years, with applications across various domains, including journalism (Liu et al., 2022; Wang et al., 2024a; Petridis et al., 2023). AI writing assistants can provide scaffolding for writers by supporting different cognitive stages of the writing process (Flower, 1981), including framing (Hui and Sprouse, 2023; Wang et al., 2024a), idea generation (Clark et al., 2018; Petridis et al., 2023) to drafting (Gero et al., 2022; Zhang et al., 2023; Long et al., 2023; Menon et al., 2024) and revision (Gero et al., 2022; Wu et al., 2019). (Lee et al., 2024) provides a comprehensive design space for intelligent and interactive writing assistants that emphasizes the importance of considering task, user, technology, interaction, and ecosystem aspects when using AI to support writers. However, using AI in creativity also has risks; overreliance (Ashktorab et al., 2021; Kim et al., 2024; Vasconcelos et al., 2023; Zhang et al., 2024) on AI can cause people to accept poor AI answers, and AI ideation can actually reduce diversity across people (Anderson et al., 2024). Broadly, AI can be seen not as a solution, but as a design material that needs to be understood by human creators (Yildirim et al., 2022; Subramonyam et al., 2021).
A major creative application of AI lies in content retargeting and transformation—editing and repurposing text, image, video, or audio for new audiences, platforms, or goals. The key is to find the fundamental message to preserve while adapting the content into the new format. Previous works have explored how to support people with content transformation tasks in long-form and short-form video creation (Truong et al., 2021; Chi et al., 2021, 2022; Wang et al., 2024b, 2022). In particular, short-form content requires restructuring narratives to fit new constraints and audience expectations (Kim and Kim, 2024).
@@ 2.3. News Reels
News reels help outlets reach younger viewers who favor TikTok and Instagram (Newman, 2022; Newman et al., 2022). Whereas traditional journalistic narratives are serious and fact-driven, news reels often use a playful tone that balances between information and entertainment, which meets expectations for social media videos (Newman, 2022).
A common way to reframe news for reels is to create a role-play. An actor/journalist creates a dialogue between two characters related to the news story and acts out the perspective of each character. For example, in a Washington Post reel, a character representing Hurricane Ian explains to a roommate character why he is “visiting” Florida and all the trouble he might cause there. Figure 1 contains an example of a news reel made in our newsroom study with a dialogue between a School District character and a Teachers Union character acting out the news of a teacher shortage. As is typical of news reels, the creator is playing both roles.
@@ 2.4. ReelFramer AI Tool
ReelFramer (Wang et al., 2024a) is a creativity support system that uses generative text and image AI to scaffold the process of creating a script and storyboard for a news reel. The user starts by inputting the news article (see Figure 2). The system uses a large language model (LLM), GPT-4, to extract locations, people, and activities from the article. The user then selects one of the three narrative framings, which contain a different balance of information and entertainment (expository dialogue, reenactment, or comedic analogy). The system then uses the LLM to suggest foundational details like characters, plot, setting, and key information, which is called the premise in scriptwriting (Batty et al., 2017). The system is iterative and built for iteration; users can accept, regenerate, or edit any of the AI generations. Once the user is happy with the script, the system provides visual framing to explore the characters’ visual design. It generates a character board that contains visual details of the costumes, props, and backgrounds to distinguish the characters using a text-to-image model, DALL-E 2. Based on the character board, it generates a storyboard that shows suggested emotions, actions, and dialogues for each character. Users can then film the reels based on these materials.
3. Student Newsroom Study
To understand the role of human creativity in the presence of an AI support tool, we present a 14-week study of how a student newsroom produces news reels using the ReelFramer AI tool.
3.1. Methods
In the student newsroom study, researchers recruited advanced students from a university journalism program that runs a TV and print journalism media outlet called AnonymizedNews. As part of their existing program, the students report local area news in print and on television with faculty guidance and approval. Before the study, AnonymizedNews had a small TikTok channel, with fewer than 50 posts and no posts getting more than 3000 views. Simulated newsroom environments are common in journalism education spaces (Steel et al., 2007) as part of practical and experiential learning with both procedural (environment) and factual (task) authenticity (Chen et al., 2001). Choosing students in a university newsroom was fairly representative of professional newsrooms because they often hire recent graduates to run social media, due to their familiarity with social media, and the demographic they aspire to reach (Newman et al., 2022).
@@ 3.1. Methods
In this simulated newsroom environment, we had one researcher acting as editor, two researchers acting as market researchers, and five students given news stories each week to convert to reels, as if they were part of a digital news desk. The news stories were ones the newsroom had previously reported. In weekly meetings, researchers observed students  (Emerson et al., 2011), and the editor interacted with the student journalists to offer the support an editor would provide (Seim, 2024) in a standard newsroom. Student journalists were regularly asked about their use or nonuse of the tool. The editor led group feedback discussions for every video and gave specific feedback for iteration. After approval by the editor, newsreels were published to the AnonymizedNews TikTok channel.
3.2. Findings: Tool Use and Disuse
@@
Overall, the student group made 26 news reels, 16 of which were published by the editor. The 10 unpublished were still being iterated at the end of the session or had been abandoned. Altogether, the 16 videos published by the team garnered over 500,000 views. Four of the five student journalists chose to use the tool for all or most of their videos. 13 of 16 published videos used ReelFramer in a significant manner to help achieve an output the editor deemed publishable.
Creators choose to use the AI tool a majority of the time—for 20 of 26 reels created. To understand what, if any, benefits the tool had, we analyzed why users used the tool and why they chose not to use the tool.
3.2.1. Reasons for Usage: Framing and Drafting Scripts
The users repeatedly mentioned the benefit of two of the steps, each supported a different cognitive process. Users mentioned not using the storyboarding features, as it was not cognitively demanding.